OpenStudy (anonymous):

Can people here answer my calc 2 questions?

OpenStudy (anonymous):

possibly, what is the problem?

OpenStudy (amistre64):

calc2? here? people with answers? .... mehbee

OpenStudy (anonymous):

Find the Integral of (sqrt(x^2-9)/x^3 You have to use trig substitutions.

OpenStudy (amistre64):

what trig subs have you tried?

OpenStudy (amistre64):

x = 3sec(t) perhaps?

OpenStudy (anonymous):

i know it's 3sec(t), i got pretty far but I got lost at the end

OpenStudy (amistre64):

x^3 = sec^3(t) dx = 3sec(t)tan(t) dt \[\int \frac{3sec(t)tan(t)}{sec^3(t)}*\sqrt{9sec^2(t)-9}\ dt\] \[\int \frac{3sec(t)tan(t)}{sec^3(t)}*3tan(t)\ dt\] \[\int 9\frac{tan^2(t)}{sec^2(t)}\ dt\] \[\int 9\frac{sin^2(t)}{cos^2(t)}*cos^2(t)\ dt\] \[9\int sin^2(t)\ dt\]

OpenStudy (amistre64):

if i did it right, the rest is simple enough

OpenStudy (amistre64):

there are a few ways to tackle a sin^2 reduction formula or indentity

OpenStudy (anonymous):

I see what you did and it looks good, but the answer in the back of the book is (1/6)sec^-1(x/3) - sqrt(x^2-9)/(2x^2) + C

OpenStudy (anonymous):

It's ridiculous, I'm not gonna waste your entire day trying to figure it out. Thanks for trying though.

OpenStudy (amistre64):

the trig sub just makes for an eaiser computation; in the end you still have to turn it back into xs

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

Thanks